# Low-Rank Graph Contrastive Learning for Node Classification

### v1

@@ 3.3 Problem Description
Noise usually exists in the input node attributes or labels of real-world graphs, which degrades the quality of the node representation obtained by common GCL encoders and affects the performance of the classifier trained on such representations. We aim to obtain node representations robust to noise in two cases, where noise is present in either the labels of $\mathcal{V_{L}}$ or in the input node attributes $\mathbf{X}$. That is, we consider either noisy label or noisy input node attributes.
The goal of LR-GCL is to learn low-rank node representations by $\mathbf{H}=g(\mathbf{X},\mathbf{A})$ such that the node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{N}$ are robust to noise in the above two cases, where $g(\cdot)$ is the LR-GCL encoder. In our work, $g$ is a two-layer GCN introduced in Section 3.2. The low-rank node representations by LR-GCL, $\mathbf{H}=\left\{\mathbf{h}_{1};\mathbf{h}_{2};\ldots;\mathbf{h}_{N}\right\}\in\mathbb{R}^{N\times d}$, are used for transductive node classification by a linear classifier. In transductive node classification, a linear transductive classifier is trained on $\mathcal{V_{L}}$, and then the classifier predicts the labels of the unlabeled test nodes in $\mathcal{V_{U}}$.
Figure 2: Illustration of the LR-GCL framework.
@@ 4 Methods
4.1 Low-Rank GCL: Low-Rank Graph Contrastive Learning
Preliminary of Prototypical GCL. The general node representation learning aims to train an encoder $g(\cdot)$, which is a two-layer Graph Convolution Neural Network (GCN) [Kipf and Welling, 2017], to generate discriminative node representations. In our work, we adopt contrastive learning to train the GCL encoder $g(\cdot)$. To perform contrastive learning, two different views, $G^{1}=(\mathbf{X}^{1},\mathbf{A}^{1})$ and $G^{2}=(\mathbf{X}^{2},\mathbf{A}^{2})$, are generated by node dropping, edge perturbation, and attribute masking. The representation of two generated views are denoted as $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$ and $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$, with $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$ being the $i$-th row of $\mathbf{H}^{1}$ and $\mathbf{H}^{2}$, respectively. It is preferred that the mutual information between $\mathbf{H}^{1}$ and $\mathbf{H}^{2}$ is maximized. For computational reason, its lower bound is usually used as the objective for contrastive learning. We use InfoNCE [Li et al., 2021] as our node-wise contrastive loss. In addition to the node-wise contrastive learning, we also adopt prototypical contrastive learning [Li et al., 2021] to capture semantic information in the node representations, which is interpreted as maximizing the mutual information between node representation and a set of estimated cluster prototypes $\left\{\mathbf{c}_{1},...,\mathbf{c}_{{K}}\right\}$. Following [Li et al., 2021, Snell et al., 2017], we use $K$-means to cluster the node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{n}$ into $K$ clusters and take the clustering centroid of the $k$-th cluster as the $k$-th prototype $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$ for all $k\in[K]$. The loss function of Prototypical GCL is comprised of two terms, $\mathcal{L}_{node}$, the loss function for node-wise contrastive learning, and $\mathcal{L}_{\textup{proto}}$, the prototypical contrastive learning loss, which are presented below:
$\small\mathcal{L}_{\textup{node}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})}{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})+\sum_{j=1}^{N}s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{j})},$
$\mathcal{L}_{\textup{proto}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)}{\sum_{k=1}^{{K}}\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)}.$ (1)
Here $s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})$ is the cosine similarity between two node representations, $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$.
@@ 4.1 Low-Rank GCL: Low-Rank Graph Contrastive Learning
LR-GCL: Low-Rank Graph Contrastive Learning. LR-GCL aims to improve the robustness and generalization capability of the node representations of Prototypical GCL by enforcing the learned feature kernel to be low-rank. The gram matrix $\mathbf{K}$ of the node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ is calculated by $\mathbf{K}=\mathbf{H}^{\top}\mathbf{H}\in\mathbb{R}^{N\times N}$. Let $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{n}$ with $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\ldots\geq\widehat{\lambda}_{\min\left\{N,d\right\}}\geq\widehat{\lambda}_{\min\left\{N,d\right\}+1}=\ldots,=0$ be the eigenvalues of $\mathbf{K}$. In order to encourage the features $\mathbf{H}$ or the gram matrix $\mathbf{H}^{\top}\mathbf{H}$ to be low-rank, we explicitly add the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}+1}\coloneqq\sum_{r=r_{0}}^{n}\widehat{\lambda}_{i}$ to the loss function of prototypical GCL. The starting rank $r_{0}<\min(n,d)$ is the rank of the features $\mathbf{H}$ we aim to obtain with the LR-GCL encoder, that is, if ${\left\|\mathbf{H}\right\|}_{r_{0}}=0$, then $\textup{rank}(\mathbf{H})=r_{0}$. Therefore, the overall loss function of LR-GCL is
$\mathcal{L}_{\textup{LR-GCL}}=\mathcal{L}_{\textup{node}}+\mathcal{L}_{\textup{proto}}+\tau{\left\|\mathbf{H}\right\|}_{r_{0}},$ (2)
@@
where $\tau>0$ is the weighting parameter for the truncated nuclear norm ${\left\|\mathbf{H}\right\|}_{r_{0}}$. We summarize the training algorithm for the LR-GCL encoder in Algorithm 1. After finishing the training, we calculate the low-rank node feature by $\mathbf{H}^{(r_{0})}=g(\mathbf{A},\mathbf{X})$.
@@
Motivation of Learning Low-Rank Features. Let ~$\tilde{}\mathbf{Y}\in\mathbb{R}^{N\times C}$ be the ground truth clean label matrix without noise. By the low frequency property illustrated in Figure 1, the projection of ~$\tilde{}\mathbf{Y}$ on the top $r$ eigenvectors of $\mathbf{K}$ with a small rank $r$, such as $r=0.2N$, covers the majority of the information in ~$\tilde{}\mathbf{Y}$. On the other hand, the projection of the label noise $\mathbf{N}$ are distributed mostly uniform across all the eigenvectors. This observation motivates low-rank features $\mathbf{H}$ or equivalently, the low-rank gram matrix $\mathbf{K}$. This is because the low-rank part of the feature matrix $\mathbf{H}$ or the gram matrix $\mathbf{K}$ covers the dominant information in the ground truth label ~$\tilde{}\mathbf{Y}$ while learning only a small portion of the label noise. Moreover, we remark that the regularization term ${\left\|\mathbf{H}\right\|}_{r_{0}}$ in the loss function (2) of LR-GCL is also theoretically motivated by the sharp upper bound for the test loss using a linear transductive classifier, presented as (A) in Theorem 4.1. A smaller ${\left\|\mathbf{H}\right\|}_{r_{0}}$ renders a smaller upper bound for the test loss, which ensures better generalization capability of the linear transductive classier to be introduced in the next subsection.
Algorithm 1 Low-Rank Graph Contrastive Learning (LR-GCL)
0:  The input attribute matrix $\mathbf{X}$, adjacency matrix $\mathbf{A}$, and the training epochs $t_{\text{max}}$.
0:  The parameters of LR-GCL encoder $g$
1:  Initialize the parameter of LR-GCL encoder $g$
2:  for $t\leftarrow 1$ to $t_{\text{max}}$ do
3:     Calculate node representations by $\mathbf{H}=g(\mathbf{X},\mathbf{A})$, generate augmented views $G^{1},G^{2}$, and calculate node representations $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$ and $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$
4:     Cluster node representations $\left\{\mathbf{h}_{i}\right\}_{i=1}^{n}$ into $K$ clusters $\left\{S_{k}\right\}_{k=1}^{K}$ with $K$-means clustering
5:     Update the prototype $\mathbf{c}_{k}$ as the centroid of $S_{k}$ by $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$ for all $k\in[K]$
6:     Calculate the eigenvalues $\left\{\lambda_{i}\right\}_{i=1}^{N}$ of the feature kernel $\mathbf{H}^{\top}\mathbf{H}$
7:     Update the parameters of LR-GCL encoder $g$ by one step of gradient descent on the loss $\mathcal{L}_{rep}$
8:  end for
9:  return The LR-GCL encoder $g$
@@ 4.2 Transductive Node Classification
In this section, we introduce a simple yet standard linear transductive node classification algorithm using the low-rank node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ produced by the LR-GCL encoder. We present strong theoretical result on the generalization bound for the test loss for our low-rank transductive algorithm with the presence of label noise.
We first give basic notations for our algorithm. Let $\mathbf{y}_{i}\in\mathbb{R}^{C}$ be the observed one-hot class label vector for node $v_{i}$ for all $i\in[N]$, and define $\mathbf{Y}\coloneqq\left[\mathbf{y}_{1};\mathbf{y}_{2};\ldots\mathbf{y}_{N}\right]\in\mathbb{R}^{N\times C}$ be the observed label matrix which may contain label noise $\mathbf{N}\in\mathbb{R}^{N\times C}$. We define $\mathbf{F}(\mathbf{W})=\mathbf{H}\mathbf{W}$ as the linear output of the transductive classier with $\mathbf{W}\in\mathbb{R}^{d\times C}$ being the weight matrix for the classifier. Our transductive classifier uses $\textup{softmax}(\mathbf{F}(\mathbf{W}))\in\mathbb{R}^{N\times C}$ for prediction of the labels of the test nodes. We train the transductive classifier by minimizing the regular cross-entropy on the labeled nodes via
@@
where KL is the KL divergence between the label $\mathbf{y}_{i}$ and the softmax of the classifier output at node $v_{i}$. We use a regular gradient descent to optimize (3) with a learning rate $\eta\in(0,\frac{1}{\widehat{\lambda}_{1}})$. We define a matrix ¯$\bar{}\mathbf{Y}(r_{0})\in\mathbb{R}^{N\times C}$ as the orthogonal projection of $\mathbf{Y}$ onto the top-$r_{0}$ eigenvectors of $\mathbf{K}$, that is, ¯$\bar{}\mathbf{Y}(r_{0})=\mathbf{U}^{(r_{0})}\left(\mathbf{U}^{(r_{0})}\right)^{\top}\mathbf{Y}$, where $\mathbf{U}^{(r_{0})}\in\mathbb{R}^{N\times r_{0}}$ is the matrix formed by the eigenvectors of $\mathbf{K}$ corresponding to the top-$r_{0}$ eigenvalues of $\mathbf{K}$. $\mathbf{W}$ is initialized by $\mathbf{W}^{(0)}=\mathbf{0}$, and at the $t$-th iteration of gradient descent for $t\geq 1$, $\mathbf{W}$ is updated by $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla_{\mathbf{W}}L(\mathbf{W})|_{\mathbf{W}=\mathbf{W}^{(t-1)}}$.
Define $\mathbf{F}(\mathbf{W},t)\coloneqq\mathbf{H}\mathbf{W}^{(t)}$ as the output of the classifier after the $t$-th iteration of gradient descent for $t\geq 1$. We have the following theoretical result on the loss of the unlabeled test nodes ${\cal V}_{{\cal U}}$ measured by the gap between $\mathbf{F}(\mathbf{W},t)$ and ¯$\bar{}\mathbf{Y}(r_{0})$ when using the low-rank feature $\mathbf{H}$ with $r_{0}\in[n]$.
@@
Let $m\geq cN$ for a constant $c\in(0,1)$, and $r_{0}\in[n]$. Assume that a set ${\cal L}$ with $\left|{\cal L}\right|=m$ is sampled uniformly without replacement from $[N]$, and a set ${\cal U}$ with $\left|{\cal U}\right|=u$ are sampled uniformly without replacement from $[N]\setminus{\cal L}$ and $m+u\leq N$. Then for every $x>0$, with probability at least $1-\exp(-x)$, after the $t$-th iteration of gradient descent for all $t\geq 1$, we have
${\cal U}_{\textup{test}}(t)\coloneqq\frac{1}{u}{\left\|\left[\mathbf{F}(\mathbf{W},t)-\mathbf{Y}\right]_{{\cal U}}\right\|}_{\textup{F}}^{2}$
$\leq\frac{c_{0}}{m}{\left\|\left(\mathbf{I}_{N}-\eta\mathbf{K}\right)^{t}\mathbf{Y}\right\|}_{\textup{F}}+c_{1}r_{0}\left(\frac{1}{u}+\frac{1}{m}\right)$
$+\left(\sqrt{\frac{{\left\|\mathbf{H}\right\|}_{r_{0}}}{u}}+\sqrt{\frac{{\left\|\mathbf{H}\right\|}_{r_{0}}}{m}}\right)+\frac{c_{2}x}{u},$ (4)

### v2

@@ 3.3 Problem Description
In real-world graph data, noise frequently appears either in the node attributes or in the labels. Such noise can significantly deteriorate the quality of node representations produced by standard GCL encoders, which in turn impairs the performance of classifiers trained on these representations. Our objective is to develop node embeddings that remain robust under two specific scenarios. The first involves noise present in the labels of $\mathcal{V_{L}}$, while the second concerns noise within the input node attributes $\mathbf{X}$. These two cases correspond to noisy labels and noisy features, respectively.
The GCL-LRR model is designed to learn node representations that are low-rank and resilient to these types of noise. The representations are defined as $\mathbf{H}=g(\mathbf{X},\mathbf{A})$, where $g(\cdot)$ denotes the GCL-LRR encoder. In this work, the encoder $g$ is implemented as a two-layer GCN, as described in Section 3.2. The output $\mathbf{H}=\left\{\mathbf{h}_{1};\mathbf{h}_{2};\ldots;\mathbf{h}_{N}\right\}\in\mathbb{R}^{N\times d}$ consists of low-rank embeddings for all nodes. These representations are then used for transductive node classification. During this process, a linear classifier is trained using the labeled node set $\mathcal{V_{L}}$, and the trained classifier is subsequently applied to predict labels for the unlabeled test nodes in $\mathcal{V_{U}}$.
@@ 4 Methods
4.1 Low-Rank GCL: Graph Contrastive Learning with Low-Rank Regularization
Preliminary of Prototypical GCL. Node representation learning focuses on training an encoder $g(\cdot)$, which is implemented using a two-layer Graph Convolutional Network (GCN) [1], to generate embeddings that are discriminative and informative. In this work, we adopt a contrastive learning framework to optimize the GCL encoder $g(\cdot)$. Two augmented graph views, denoted as $G^{1}=(\mathbf{X}^{1},\mathbf{A}^{1})$ and $G^{2}=(\mathbf{X}^{2},\mathbf{A}^{2})$, are created by applying transformations such as node dropping, edge perturbation, and attribute masking. The resulting node representations are given by $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$ and $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$, where $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$ refer to the representations of the $i$-th node in the respective views.
The objective is to enhance the mutual information between $\mathbf{H}^{1}$ and $\mathbf{H}^{2}$, and this is achieved by optimizing a lower bound on mutual information through the InfoNCE loss [41], which serves as the node-level contrastive objective. In order to encode more semantic structure in the learned representations, we also incorporate prototypical contrastive learning [41], which encourages alignment between node embeddings and a set of cluster prototypes $\left\{\mathbf{c}_{1},...,\mathbf{c}_{{K}}\right\}$. Following the approaches described in [41, 42], we perform $K$-means clustering on the set $\left\{\mathbf{h}_{i}\right\}_{i=1}^{N}$ to obtain the $K$ clusters $\left\{S_{k}\right\}_{k=1}^{K}$. Each cluster prototype $\mathbf{c}_{k}$ is computed as $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$ for every $k$ in $[K]$. The overall objective for Prototypical GCL consists of two components, $\mathcal{L}_{\textup{node}}$, which corresponds to node-level contrastive loss, and $\mathcal{L}_{\textup{proto}}$, which accounts for the prototype-based contrastive learning. $\mathcal{L}_{\textup{node}}$ and $\mathcal{L}_{\textup{proto}}$ are computed as
$\mathcal{L}_{\textup{node}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})}{s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})+\sum_{j=1}^{N}s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{j})},$
$\mathcal{L}_{\textup{proto}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)}{\sum_{k=1}^{{K}}\exp(\mathbf{H}_{i}\cdot\mathbf{c}_{k}/\tau)},$
where $s(\mathbf{H}^{1}_{i},\mathbf{H}^{2}_{i})$ is the cosine similarity between $\mathbf{H}^{1}_{i}$ and $\mathbf{H}^{2}_{i}$.
@@ 4.1 Low-Rank GCL: Graph Contrastive Learning with Low-Rank Regularization
GCL-LRR: Graph Contrastive Learning with Low-Rank Regularization. GCL-LRR enhances the robustness and generalization of node representations derived from Prototypical GCL by promoting a low-rank learned feature kernel. The gram matrix $\mathbf{K}$ of the node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ is calculated by $\mathbf{K}=\mathbf{H}^{\top}\mathbf{H}\in\mathbb{R}^{N\times N}$. Let $\left\{\widehat{\lambda}_{i}\right\}_{i=1}^{N}$ with $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\ldots\geq\widehat{\lambda}_{\min\left\{N,d\right\}}\geq\widehat{\lambda}_{\min\left\{N,d\right\}+1}=\ldots,=0$ be the eigenvalues of $\mathbf{K}$. In order to encourage the features $\mathbf{H}$ or the gram matrix $\mathbf{H}^{\top}\mathbf{H}$ to be low-rank, we explicitly add the TNN ${\left\|\mathbf{K}\right\|}_{r_{0}}\coloneqq\sum_{r=r_{0}+1}^{N}\widehat{\lambda}_{i}$ to the loss function of prototypical GCL. The starting rank $r_{0}<\min(N,d)$ is the rank of the gram matrix of the features we aim to obtain with the GCL-LRR encoder, that is, if ${\left\|\mathbf{K}\right\|}_{r_{0}}=0$, then $\textup{rank}(\mathbf{K})=r_{0}$. Therefore, the overall loss function of GCL-LRR is
$\mathcal{L}_{\textup{GCL-LRR}}=\mathcal{L}_{\textup{node}}+\mathcal{L}_{\textup{proto}}+\tau{\left\|\mathbf{K}\right\|}_{r_{0}},$ (1)
@@
where $\tau>0$ is the weighting parameter for the TNN ${\left\|\mathbf{K}\right\|}_{r_{0}}$. We summarize the training algorithm for the GCL-LRR encoder in Algorithm 1. After the training is completed, we compute the low-rank node feature by $\mathbf{H}=g(\mathbf{A},\mathbf{X})$. Algorithm 1 outlines the training procedure for Graph Contrastive Learning with Low-Rank Regularization (GCL-LRR). The characteristics of the datasets employed in our experimental evaluations are summarized in Table 2.
Algorithm 1 Graph Contrastive Learning with Low-Rank Regularization (GCL-LRR)
1:Attribute matrix $\mathbf{X}$, adjacency matrix $\mathbf{A}$, training epochs $t_{\text{max}}$
2:Parameters of GCL-LRR encoder $g$
3:Initialize parameters of encoder $g$
4:for $t=1$ to $t_{\text{max}}$ do
5:     Compute node representations: $\mathbf{H}=g(\mathbf{X},\mathbf{A})$
6:     Generate augmented views $G^{1},G^{2}$ with $\mathbf{X}^{1},\mathbf{A}^{1}$ and $\mathbf{X}^{2},\mathbf{A}^{2}$
7:     Compute $\mathbf{H}^{1}=g(\mathbf{X}^{1},\mathbf{A}^{1})$, $\mathbf{H}^{2}=g(\mathbf{X}^{2},\mathbf{A}^{2})$
8:     Cluster $\{\mathbf{h}_{i}\}_{i=1}^{N}$ into $K$ clusters $\{S_{k}\}_{k=1}^{K}$ using $K$-means
9:     for $k=1$ to $K$ do
10:         Compute prototype: $\mathbf{c}_{k}=\frac{1}{|S_{k}|}\sum_{\mathbf{h}_{i}\in S_{k}}\mathbf{h}_{i}$
11:     end for
12:     Compute eigenvalues $\{\lambda_{i}\}_{i=1}^{N}$ of kernel matrix $\mathbf{K}=\mathbf{H}^{\top}\mathbf{H}$
13:     Update parameters of $g$ using gradient descent on $\mathcal{L}_{\textup{GCL-LRR}}$
14:end for
15:return Encoder $g$
Figure 1: Eigen-projection (first row) and signal concentration ratio (second row) on Cora, Citeseer, and Pubmed. The study in this figure is performed for asymmetric label noise with a noise level of $60\%$. By the rank $r=0.2\min\left\{N,d\right\}$, the signal concentration ratio of ~$\tilde{}\mathbf{Y}$ for Cora, Citeseer, and Pubmed are $0.844$, $0.809$, and $0.784$ respectively. Figure 2 in Section 5.8 further illustrates the eigen-projection and signal concentration ratio on more datasets.
@@
Motivation of Learning Low-Rank Features. We investigate how the information from the ground-truth clean labels and the label noise is distributed across different eigenvectors of the feature gram matrix $\mathbf{K}$ through an eigen-projection analysis. Let ~$\tilde{}\mathbf{Y}\in\mathbb{R}^{N\times C}$ denote the clean label matrix without noise. We begin by computing the eigenvectors $\mathbf{U}$ of the gram matrix $\mathbf{K}=\mathbf{H}\mathbf{H}^{\top}$. The eigen-projection score for the $r$-th eigenvector is then given by $p_{r}=\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}^{(r)}}^{\top}\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}/{\left\|\tilde{}\mathbf{Y}^{(c)}\right\|}_{2}^{2}$ for $r\in[N]$, where $C$ is the number of classes, and ~$\tilde{}\mathbf{Y}\in{0,1}^{N\times C}$ consists of one-hot encoded clean labels. Here ~$\tilde{}\mathbf{Y}^{(c)}$ refers to the $c$-th column of ~$\tilde{}\mathbf{Y}$. We define $\mathbf{p}=\left[p_{1},\ldots,p_{N}\right]\in\mathbb{R}^{N}$ as the vector of projection values. In the presence of label noise $\mathbf{N}\in\mathbb{R}^{N\times C}$, the observed label matrix becomes $\mathbf{Y}=\tilde{}\mathbf{Y}+\mathbf{N}$. The projection value $p_{r}$ quantifies the proportion of the signal aligned with the $r$-th eigenvector of $\mathbf{K}$, while the signal concentration ratio at rank $r$ of the ground truth class label is defined as ${\left\|\mathbf{p}^{(1:r)}\right\|}_{1}$, representing the cumulative contribution of the top $r$ eigenvectors. Similarly, the signal concentration ratio at rank $r$ of the noise is defined as $\frac{1}{C}\sum_{c=1}^{C}{\left\|{\mathbf{U}^{(r)}}^{\top}\mathbf{N}^{(c)}\right\|}_{2}^{2}/{\left\|\mathbf{N}^{(c)}\right\|}_{2}^{2}\in\mathbb{R}^{N}$. Empirical results, shown as red curves in Figure 1, indicate that the clean label signals are primarily concentrated on the leading eigenvectors of $\mathbf{K}$. In contrast, the projection of the label noise appears more uniformly distributed across all eigenvectors, as demonstrated by the blue curves in the same figure. The above observation motivates low-rank features $\mathbf{H}$ or, equivalently, the low-rank gram matrix $\mathbf{K}$. This is because the low-rank part of the feature matrix $\mathbf{H}$ or the gram matrix $\mathbf{K}$ covers the dominant information in the ground truth label ~$\tilde{}\mathbf{Y}$ while learning only a small portion of the label noise. We refer to such property as the Low Frequency Property (LFP), which has been widely studied in deep learning [66, 67, 68, 69, 70]. Moreover, we remark that the regularization term ${\left\|\mathbf{K}\right\|}_{r_{0}}$ in the loss function (1) of GCL-LRR is also theoretically motivated by the sharp upper bound for the test loss using a linear transductive classifier, presented as (3) in Theorem 4.1. A smaller ${\left\|\mathbf{K}\right\|}_{r_{0}}$ renders a smaller upper bound for the test loss, which ensures better generalization capability of the linear transductive classifier to be introduced in the next subsection.
@@ 4.2 Transductive Node Classification
In this section, we present a standard yet straightforward linear transductive node classification method based on the low-rank node representations $\mathbf{H}\in\mathbb{R}^{N\times d}$ obtained from the GCL-LRR encoder. Furthermore, we establish a novel generalization bound on the test loss of this transductive classifier in the presence of label noise.
We begin by introducing the fundamental notations for the proposed algorithm. For each node $v_{i}$ where $i\in[N]$, let $\mathbf{y}_{i}\in\mathbb{R}^{C}$ denote its observed one-hot class label vector. We then define the full observed label matrix as $\mathbf{Y}\coloneqq\left[\mathbf{y}_{1};\mathbf{y}_{2};\ldots;\mathbf{y}_{N}\right]\in\mathbb{R}^{N\times C}$, which may include label noise, modeled as $\mathbf{N}\in\mathbb{R}^{N\times C}$. The classifier’s linear prediction is defined by $\mathbf{F}(\mathbf{W})=\mathbf{H}\mathbf{W}$, where $\mathbf{W}\in\mathbb{R}^{d\times C}$ is the learnable weight matrix. Final predictions are made using the softmax transformation $\textup{softmax}(\mathbf{F}(\mathbf{W}))\in\mathbb{R}^{N\times C}$ to estimate class probabilities for the test nodes. We train the transductive classifier by minimizing the regular cross-entropy on the labeled nodes through the following optimization problem,
@@
where KL is the KL divergence between the label $\mathbf{y}_{i}$ and the softmax of the classifier output at node $v_{i}$. We use a regular gradient descent to optimize (2) with a learning rate $\eta\in(0,\frac{1}{\widehat{\lambda}_{1}})$. $\mathbf{W}$ is initialized by $\mathbf{W}^{(0)}=\mathbf{0}$, and at the $t$-th iteration of gradient descent for $t\geq 1$, $\mathbf{W}$ is updated by $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla_{\mathbf{W}}L(\mathbf{W})|_{\mathbf{W}=\mathbf{W}^{(t-1)}}$. We define $\mathbf{F}(\mathbf{W},t)\coloneqq\mathbf{H}\mathbf{W}^{(t)}$ as the output of the classifier after the $t$-th iteration of gradient descent for $t\geq 1$. We have the following theoretical result, Theorem 4.1, on the Mean Squared Error (MSE) loss of the unlabeled test nodes ${\cal V}_{{\cal U}}$ measured by the gap between $\left[\mathbf{F}(\mathbf{W},t)\right]_{{\cal U}}$ and [~$\left[\tilde{}\mathbf{Y}\right]_{{\cal U}}$ when using the low-rank feature $\mathbf{H}$ with $r_{0}\in[N]$, which is the generalization error bound for the linear transductive classifier using $\mathbf{F}(\mathbf{W})=\mathbf{H}\mathbf{W}$ to predict the labels of the unlabeled nodes. Similar to existing works such as [71] that use the Mean Squared Error (MSE) to analyze the optimization and the generalization of GNNs, we employ the MSE loss to provide the generalization error of the node classifier in the following theorem. It is remarked that the MSE loss is necessary for the generalization analysis of transductive learning using the transductive local Rademacher complexity [72, 73].
@@
Let $m\geq cN$ for a constant $c\in(0,1)$, and $r_{0}\in[N]$. Assume that a set ${\cal L}$ with $\left|{\cal L}\right|=m$ is sampled uniformly without replacement from $[N]$, and the remaining nodes ${\cal V}_{U}={\cal V}\setminus{\cal V}_{L}$ are the test nodes. Then for every $x>0$, with probability at least $1-\exp(-x)$, after the $t$-th iteration of gradient descent for all $t\geq 1$, we have